### v1

# Enhancing Wide-Angle Image Using Narrow-Angle View of the Same Scene

Abstract
A common dilemma while photographing a scene is whether to capture it in wider angle, allowing more of the scene to be covered but in lesser details or to click in narrow angle that captures better details but leaves out portions of the scene. We propose a novel method in this paper that infuses wider shots with finer quality details that is usually associated with an image captured by the primary lens by capturing the same scene using both narrow and wide field of view (FoV) lenses. We do so by training a GAN-based model to learn to extract the visual quality parameters from a narrow angle shot and to transfer these to the corresponding wide-angle image of the scene. We have mentioned in details the proposed technique to isolate the visual essence of an image and to transfer it into another image. We have also elaborately discussed our implementation details and have presented the results of evaluation over several benchmark datasets and comparisons with contemporary advancements in the field.
keywords:
Ultra-wide Image , Super Resolution , Generative Adversarial Network
[…]
[auth1]organization=Department of Computer Science and Engineering, University of Dhaka, city=Dhaka
[…]
[auth2]organization=Department of Computer Science and Engineering, Jahangirnagar University, city=Dhaka
[…]
[auth3]organization=Department of Computer Science and Engineering, University of Dhaka, city=Dhaka
1 Introduction
Ultra-wide photography is a popular photographic trend in today’s world. This form of image-taking, though accommodating a greater field of view and thus offering a unique aesthetic, degrades the finer details better retained by the primary lens. This trade-off between the field of view covered and the quality of the finer details retained by the captured image places the users in an unwanted dilemma where one has to compromise the field of view for the details or vice-versa. Our primary goal in this paper is to enable the camera users to capture wide-angle images with quality that is identical to the narrow-angle shots while covering the original wide field of view.
The fundamental idea of capturing wide-angle images that retain the finer details associated with narrow-angle shots is to capture the scene of interest with two different lenses, both narrow and wide. The expectation here is that the narrow-angle shot provides the expected resolution while the wide-angle one carries the contents of the scene not covered in the narrow-angle counterpart. The idea would then be to render the contents captured by the wide-angle lens in the resolution that is represented by the narrow-angle counterpart. As it has been shown by the paper on Neural Style Transfer [6], the style and content of an image are separable and the stylistic representation of an image can be transferred into another image. We have leveraged this finding to transfer the resolution qualities of the narrow-angle shot into the wide-angle counterpart. Furthermore, we have implemented a pipeline that incrementally transfers the qualitative representations from the narrow-angle shot to the immediate wide-angle shot and repeats the process until the widest available image (often 0.5x or ultra-wide image) is reached. This involves capturing the scene of interest using all available lenses and incrementally applying the finest available quality (from the image with the narrowest field of view) through the stack of available images up to the image with the widest field of view.
In this paper, we propose a GAN-based architecture where the generator network produces realistic-looking enhanced versions of the wide FoV image while the discriminator component attempts to identify generated high-resolution results from the ground truth. In the first step, the spatial resolution of the original wide FoV image is upscaled using bicubic interpolation so that more details can be physically accommodated. Once upscaled, both the enhanced wide (from now onwards to be referred to as the wide FoV) and the narrow FoV images are partitioned into smaller patches so that visually similar patches may be utilized to assist the process of resolution transfer, inspired by the paper on Vision Transformer [11]. For the purpose of resolution transfer, we propose a modified version of the Gram matrix representation of an image, as used in the paper on Neural Style Transfer. Upon transferring the resolution from the narrow FoV patches into that from the wide FoV image, the synthetic image is passed through the discriminator setup that refines the seams between the tiles and the overall perceptual fidelity of the image, leveraging the task-specific loss functions we have designed. Finally, we perform this routine of resolution transfer throughout the stack of all available lenses to attain the highest possible resolution for the widest available field of view. Our contributions can thus be summarized into the following points,
  • 1.
    An end-to-end pipeline capable of transferring finer quality details to the wide-angle image from the corresponding narrow-angle image
  • 2.
    The novel concept of transferring visual quality across the stack of all available lenses
  • 3.
    An innovative modification of the Gram matrix that is able to capture the visual essence of an image
2 Related Works
Enhancing image quality has remained a central pursuit in computer vision. Earlier methods were primarily centered around super-resolution or upscaling techniques [1], which aimed to reconstruct high-resolution images from their low-resolution counterparts. While effective in producing sharper visuals, these models often introduced artifacts like oversmoothing, leading to a loss in texture and fine details. Later, generative adversarial models such as SRGAN [3], ESRGAN [2], and Real-ESRGAN [4] improved perceptual realism by emphasizing texture fidelity and producing visually convincing outputs. However, these models still struggle with preserving structural consistency, especially in scenes where the narrow and wide FOV images exhibit differences in perspective or content.
Another stream of relevant research comes from neural style transfer. Initially developed for artistic applications [6], these models demonstrated that image attributes could be transferred from one image to another. Subsequent work improved efficiency using feed-forward networks [20] and adaptive normalization techniques [8], broadening their use cases. However, transferring quality between narrow and wide FOV images introduces unique challenges not fully addressed by style transfer methods, as the wide FOV often contains additional contextual elements that require accurate integration.
Recent advancements have introduced lightweight and efficient architectures tailored for real-time applications. LatticeNet [24] employs a lattice block structure inspired by filter banks, enabling adaptive feature fusion while maintaining low computational complexity. This design facilitates the reconstruction of high-quality images without the overhead associated with deeper networks.
Building upon the success of attention mechanisms, the Efficient Long-Range Attention Network (ELAN) [25] integrates shift convolutions with group-wise multi-scale self-attention modules. This combination captures both local and long-range dependencies, enhancing the model’s ability to reconstruct fine details across varying scales. ELAN’s architecture achieves a balance between performance and efficiency, making it suitable for applications requiring real-time processing. Further exploring attention-based models, DLGSANet [23] introduces a hybrid approach combining dynamic local and sparse global self-attention mechanisms. The Multi-Head Dynamic Local Self-Attention (MHDLSA) module effectively captures local features, while the Sparse Global Self-Attention (SparseGSA) module focuses on the most relevant global information. This dual attention strategy enhances the model’s capacity to reconstruct high-resolution images with reduced computational demands. In parallel, the Deep Multi-level Feature Fusion Network (DMFFN) [26] addresses common challenges in super-resolution tasks, such as blurred edges and redundant structures. By employing two-level nested residual blocks with asymmetric structures, DMFFN efficiently fuses multi-level features, leading to improved edge preservation and faster convergence during training.
These contemporary models demonstrate significant progress in balancing performance and efficiency, particularly in scenarios requiring the enhancement of wide FOV images by leveraging information from narrow FOV counterparts. Their architectures are designed to maintain structural consistency and fine details, addressing the limitations observed in earlier methods.
Moreover, the integration of perceptual metrics such as LPIPS [16] has become increasingly prevalent, offering a more nuanced assessment of visual similarity that aligns closely with human perception. Incorporating human-in-the-loop evaluations further ensures that the enhanced images meet real-world expectations for quality and accuracy.
In this work, building upon these advancements, we aim to improve wide FOV image quality by transferring visual fidelity from narrow FOV versions of the same scene. By combining insights from generative models, reference-based super-resolution, attention mechanisms, and perceptual consistency techniques, we target high-quality, structurally accurate image synthesis suited for immersive visual applications.
3 Method
In this section, we have described in detail the adopted methodologies and architecture of the model used to transfer high-quality details from the narrow-angle image into the wide-angle counterpart.
Figure 1: Model Architecture
In the case of training the model, we consider the original image as the wide FoV ground truth and extract the central portion to serve as the high-definition narrow FoV. The original wide FoV is then downsampled to simulate the captured wide FoV image.
3.1 Content Reconstruction
Our primary goal is to reconstruct the contents of the wide-angle image applying the visual qualities of the narrow-angle counterpart. By visual qualities we indicate the parameters that impact the perceptual cues which are color, brightness, contrast, and sharpness. For reconstructing the contents of the wide-angle image, we utilize the pixel-wise content loss. We predict the pixel entries for a white-noise image such that the following content loss is minimized.
$L^{(l)}_{content}(p,q)=\frac{1}{2}\sum_{i,j}(Q^{l}_{ij}-P^{l}_{ij})^{2}$
Here $p$ and $q$ are the original and white-noise images respectively and $P^{l}_{ij}$ and $Q^{l}_{ij}$ are the feature representations in the $l^{th}$ layer for $i^{th}$ filter at the position $j$.
3.2 Visual Quality Reconstruction
As mentioned earlier, the visual qualities are obtained from the narrow-angle image. The qualities are extracted from the image using the gram matrix approach. We have divided the image into a fixed number of patches and computed Gram matrices for each patch. The motivation to adopt this approach comes from the intuition that visually similar patches may exist between the narrow and the wide-angle images. Obtaining visual qualities from the patch of narrow-angle image that visually aligns with a patch of the wide-angle image would produce better results. Hence, we divide the entire narrow and wide-angle images into a predefined number of patches. For each patch of the wide-angle image, the patch from the narrow-angle image that is visually most similar is selected. The visual similarity score between patches is determined using a perceptual loss function described in the following section. Patches are taken to be visually similar if the similarity score exceeds a threshold. In case no patch from the narrow-angle image is found to exceed the threshold, a gram matrix representation of the entire narrow-angle image is used as a backup.
Figure 2 illustrates partitioning the image into fixed-sized patches. Images partitioned into fixed-sized patches just for illustrative purposes do not indicate the actual number of patches or their proportions used in the implementation. The patches are bound by black lines. Arrows and bounding boxes of different colors denote patches in the narrow FoV image that are visually similar to certain patches from the wide FoV image.
Figure 2: Image partitioning and patch similarity between narrow and wide FoV images.
3.2.1 Dividing into Patches
The narrow and wide-angle images are first divided into patches of dimension $128\times 128$. This particular size for the patches has been chosen to create a balance between structural information contained within an image and the computational cost incurred to process the patches. The idea of dividing the images into patches also takes inspiration from the paper on Vision Transformer [11].
3.2.2 Patch-wise Similarity
We use perceptual loss [17] to quantify the visual similarity between patches from narrow and wide-angle images. We use the mid layers of a VGG19 network to numerically determine the perceptual loss between the pair of patches. Following function is used to determine the loss,
$L_{feature}(p_{1},p_{2})=\frac{1}{C_{j}*H_{j}*W_{j}}||\phi_{j}(p_{2})-\phi_{j}(p_{1})||_{2}^{2}$
Here $p_{1}$ and $p_{2}$ are the patches from wide and narrow-angle images, $\phi_{j}(p_{1})$ and $\phi_{j}(p_{2})$ denote the feature map obtained from the $j^{th}$ layer, and $C_{j},H_{j}$ and $W_{j}$ are the number of channels, height, and width of the patches.
3.3 Generator Architecture
Figure 3: Generator Architecture
We utilize a ResNet [27] based generator with residual blocks. The input layer is followed by a collection of residual blocks to facilitate earlier layers in learning complex functions. The residual block relays the data into the upsampling block that generates the enhanced patch and passes through a final convolution layer to adjust the channel dimension.
3.3.1 Feature Map Construction
We consider the following parameters to effectively represent the visual qualities of an image (patch of an image in our case); $color,contrast,brightness,sharpness$. Considering each of these parameters as distinct features of a patch, an enhanced feature map representation of the patch is computed. The enhanced feature map is an augmentation of the individual features extracted from the patch and is formulated as follows,
$\hat{F_{l}}=[F_{l},F_{l_{color}},F_{l_{brightness}},F_{l_{contrast}},F_{l_{sharpness}}]$
Where $F_{l}\epsilon\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ is the feature map extracted using a VGG backbone network. Rest of the features are computed as follows,
  1. 1.
    Color: An average of the RGB channels forms an embedding that represents the overall intensities. Computed as,
    $F_{l_{color}}=\frac{1}{H_{l}W_{l}}\sum^{H_{l}}_{i=1}\sum^{W_{l}}_{j=1}F_{l}$
  2. 2.
    Brightness: Mean pixel intensities over all pixels are taken to indicate the perceived luminance of the patch. Computed as,
    $F_{l_{brightness}}=\frac{1}{C_{l}H_{l}W_{l}}\sum^{C_{l}}_{i=1}\sum^{H_{l}}_{i=1}\sum^{W_{l}}_{k=1}F_{l}$
  3. 3.
    Contrast: Variance of the pixel intensities is computed to represent the variation of intensities, that is the contrast, within the patch. Calculated as,
    $F_{l_{contrast}}=\sqrt{\frac{1}{C_{l}H_{l}W_{l}}\sum^{C_{l}}_{i=1}\sum^{H_{l}}_{i=1}\sum^{W_{l}}_{k=1}(F_{l}-\mu_{l})^{2}}$
    here $\mu_{l}$ is the average intensity.
  4. 4.
    Sharpness: Sharpness is captured using the Laplacian filter. The Laplacian filter is multiplied to the feature map as follows,
    $F_{l_{sharpness}}=F_{l}*\bigtriangledown^{2}$
3.3.2 Gram Matrix Computation
Once the feature map is constructed, we can use this representation to find the gram matrix representation of the patch of the image for the current layer. The computation can be performed as follows,
$G_{l}=\hat{F}_{l}\hat{F}_{l}^{\top}$
The gram matrix is the product of the enhanced feature map to its transpose, essentially capturing the pair-wise correlation between features.
3.3.3 Loss Calculation
Visual quality loss is defined as the squared distance between the gram matrix representation of the source and target images. In our case, the source is the patch from the narrow FoV image while the target is a white-noise image which is eventually going to ”be” the resultant image by minimizing the loss defined above. The numerical representation of the loss is,
$L^{(l)}_{visual}=\sum_{l}w_{l}||G_{l_{q}}-G_{l_{p}}||^{2}$
Here $G_{l_{q}}$ and $G_{l_{p}}$ are the gram matrix representations of the patches from the white-noise and the narrow FoV images respectively. Mentionable that in case no particular match is found for a certain wide FoV patch in the narrow FoV, $G_{l_{p}}$ would denote the gram matrix representation of the entire narrow FoV image.
Hence the total generator loss is quantified as,
$L^{(l)}_{G}=L^{(l)}_{content}+L^{(l)}_{visual}$
3.4 Discriminator Loss
For the generated output image to be realistic and visually plausible, the results generated by the deconvolution network is passed through a discriminator component that aims to differentiate synthetic images from the real ones. We have designed two loss functions to assist the discriminator in this task. The losses are described as follows,
3.4.1 Seam Consistency Loss
This loss is aimed at eliminating the seams between the patches in the synthetic image by minimizing the perceptual difference at the boundaries of the adjacent patches. For adjacent patches $P_{i}$ and $P_{j}$, we take neighboring regions $R_{i}$ and $R_{j}$ which are $w$ pixels wide and compute the loss as,
$L^{(l)}_{seam}=\frac{1}{w}\sum||\phi_{l}(R_{j})-\phi_{l}(R_{i})||^{2}$
3.4.2 Perceptual Loss against Ground Truth
To ensure the essence of the narrow FoV has been plausibly transferred to the synthetic image, the output is perceptually compared to the original wide FoV ground truth. We use the same VGG19 activation maps to find the perceptual difference $L^{(l)}_{perceptual}$ between the generated wide FoV image and the original image from the dataset.
Total discriminator loss is defined as follows,
$L^{(l)}_{D}=L^{(l)}_{seam}+L^{(l)}_{perceptual}$
3.5 Cascading Lens Stack
The above procedure is incrementally applied across the available lens array to render the widest possible image with a resolution comparable to the narrowest available shot of the scene. For example, if a camera has lenses with the following zoom capacities: $5x,3x,1x,0.5x$; visual qualities are iteratively transferred from the $5x$ to $3x$, then from $3x$ to $1x$ and so on until the widest available shot has been reached.
4 Experimentation
4.1 Datasets
We have used two datasets to ensure robust training of our proposed model.
  • 1.
    DIV8K [19]: High-resolution images widely used for super-resolution tasks. More than 1500 photos with a maximum resolution of 8K are included in the dataset. It covers a wide range of scene contents.
  • 2.
    Landscapes Dataset [18] (LHQ 1024×1024): With 90,000 high-quality images of 1024 x 1024 resolution, it includes a diverse range of urban and natural environments. These high-quality landscape images with rich textures and details challenge the method’s ability to handle complex natural scenes.
4.2 Generator
As stated earlier, the generator network takes in $128\times 128\times 3$ patches and performs an initial convolution followed by a ReLU layer. The feature map is then fed into an array of $64$ residual blocks. Residual blocks assist in preventing the gradient from dying out during back propagation. This is followed by $2$ upsampling blocks, responsible for physically upscaling the patch using Pixel Shuffle layers. The upsampled patches passed through the output convolution block to adjust the channel dimension and come out as enhanced patches with a dimension of $256\times 256\times 3$.
4.3 Discriminator
We use a ResNet18 based discriminator network that takes as input a $1024\times 1024\times 3$ image formed by stitching the enhanced patches generated by the generator using Poisson Blending. The image undergoes a series of $6$ residual blocks followed by a layer of adaptive pooling and a fully connected layer outputting a single scaler denoting the realness of the input image.
4.4 Simulating Narrow and Wide FoVs
A high-quality narrow FoV and two wide FoV (one low-quality and one high-quality) images are required for training. High-quality wide images are used to calculate the loss, whereas low-quality wide and high-quality narrow images are used as model inputs. The original images from the dataset are treated as high-quality wide FoV images as well as the ground truth. To simulate high-quality narrow and low-quality wide-angle images from single-image datasets we have used the following techniques.
High-quality narrow FoV simulation: For each image in the dataset, we extract a high-resolution central region with a zoom factor of $\frac{5}{3}$. This region is then embedded into a larger blank canvas, preserving the original central region’s quality. The resulting image replicates the characteristics of a narrow FoV while maintaining the high quality of the original.
Low-quality wide FoV simulation: For the DIV8K dataset with 8K resolution, we simulated low-quality wide FoV images by first downsampling each image to 4096 × 4096 and then upsampling it back to 8K. Similarly, for the Landscapes dataset (LHQ 1024 × 1024), we downsampled each image to 512 × 512 and then upsampled it back to 1024 × 1024. This intentional degradation process mimics low-quality imaging conditions while preserving the broader scene context, allowing us to test our methods under challenging visual conditions. The description of workflow provided in this paper is in light of the Landscape dataset, that is, assuming the original image to be 1024 x 1024. The image reconstruction pipeline has been shown in the following Figure-4.
Figure 4: Enhanced image reconstruction pipeline
4.5 Training Details
Key hyperparameters for training and evaluation included a patch size of 128 × 128, Adam optimizer with a learning rate of 0.0001 with exponential decay every 10 epochs, a batch size of 16, perceptual loss layers (Conv2_2, Conv3_2, Conv4_2 of VGG19), and a threshold for patch similarity empirically set to 0.7.
The generator network was pre-trained for 50 epochs using content and visual losses before incorporating adversarial loss. The full generator-discriminator framework was then trained for an additional 200 epochs, totaling 400K iterations. The generator consists of 16 residual blocks, each with two 3×3 convolutional layers, batch normalization, and parametric ReLU activation, allowing effective feature extraction while keeping training stable. To enhance high-frequency details, two upsampling layers using pixel shuffle progressively increase the resolution. The final 3×3 convolutional layer reconstructs the high-resolution image.
The convergence behavior of the visual quality loss for a single patch in the generator network is shown in Figure 5. It highlights the effectiveness of our proposed enhanced Gram matrix, which causes an earlier convergence by approximately 30K iterations compared to the traditional approach. This earlier stabilization suggests that our method improves the efficiency of optimization, reducing the training time required while maintaining reconstruction quality.
Figure 5: Faster convergence of visual quality loss with the enhanced Gram matrix.
To further stabilize the training and prevent mode collapse, spectral normalization was applied to the discriminator while gradient clipping controlled exploding gradients. The model was optimized using a combination of content loss (L1), perceptual loss (VGG-based), and adversarial loss, ensuring a strong balance between detail accuracy and perceptual quality.
4.6 Experiment Results
Table 1 presents the results of our experiment, where a narrow and a low-resolution wide image are used as input. These images are processed through our model to generate the final output.
Table 2 illustrates the cascading effect of the outputs derived from images captured at different zoom levels. In the first row, the input images consist of a 2x (telephoto) and a 1x (standard) zoom, which collectively produce a high-resolution 1x (standard) output. The second row takes the output from the previous row as one input and combines it with a low-resolution 0.5x (wide) image, ultimately generating a high-resolution 0.5x (wide) image. This sequence effectively simulates how our method can be applied in a cascading manner using all available zoom lenses.
Narrow FoV Input wide FoV Our result
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 1: Experimental results showing the output generated by our model when processing narrow and wide input images.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
2x Zoom 1x Zoom 1x Output
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
1x Output 0.5x Zoom 0.5x Output
Table 2: Experimental results demonstrating the cascading effect of multi-zoom image upscaling.
5 Comparison with Prior Arts
We have evaluated our proposed methods on PSNR, SSIM and LPIPS[16] metrics over Set5[12], Set14[13] and BSD100[14] datasets. For PSNR and SSIM, we have compared our results with existing technologies such as ESRGAN[2], LatticeNet[24], ELAN[25], DLGSANet[23] and DMFFN[26] in Table 3 and Table 4.
Dataset ESRGAN LatticeNet ELAN DLGSANet DMFFN Ours
Set5 32.73 32.18 32.43 32.46 32.41 32.80
Set14 28.99 28.61 28.78 28.79 28.74 30.37
BSD100 25.16 27.57 27.69 27.70 27.74 26.58
Table 3: PSNR Comparison (highest marked in red)
Dataset ESRGAN LatticeNet ELAN DLGSANet DMFFN Ours
Set5 0.9011 0.8943 0.8975 0.8993 0.8981 0.9118
Set14 0.7917 0.7812 0.7858 0.7871 0.7874 0.7806
BSD100 0.7455 0.7355 0.7406 0.7415 0.7421 0.8261
Table 4: SSIM Comparison (highest marked in red)
We have presented our comparison with bicubic interpolation and Multiscale GAN [15] on the LPIPS metric in Table 5. It is to be noted that LPIPS values are within the range $[0,1]$, with a lower value denoting the generated image is perceptually more aligned with the ground truth.
Dataset bicubic Multiscale GAN Ours
Set5 0.5657 0.2823 0.0795
Set14 0.6114 0.3717 0.2032
BSD100 0.5726 0.4112 0.3361
Table 5: LPIPS Comparison (best marked in red)
Figure 6 presents a comparative analysis of our method against several prior super-resolution models, including bicubic interpolation, LatticeNet [24], ELAN [25], and DLGSANet [23]. The leftmost image represents the original ground truth, with the red bounding box highlighting the region of interest for focused evaluation. The comparative results illustrate variations in texture reconstruction, detail preservation, and structural fidelity across different approaches. Notably, our method, displayed in the bottom-right corner, achieves superior sharpness and consistency, effectively reducing blurring and artifacts observed in other models.
Figure 6: Visual comparison of super-resolution models, including bicubic interpolation, LatticeNet, ELAN, and DLGSANet. The red bounding box highlights the region of interest, with our method (bottom-right) achieving superior detail preservation and structural fidelity
6 Conclusion
We proposed a GAN-based approach to enhance wide-angle images by transferring fine details from a narrow-angle shot. Our method balances coverage and detail through multi-zoom image fusion, leveraging perceptual losses and an improved Gram matrix representation. The cascading lens stack further extends its application to multi-lens imaging systems.
Extensive experiments showed that our approach outperforms existing super-resolution and image synthesis methods, achieving superior PSNR, SSIM, and LPIPS scores while reducing artifacts. This opens new possibilities for computational photography, smartphone imaging, and digital zoom. Future work will focus on improving performance under extreme lighting, dynamic scenes, and real-time constraints, with potential enhancements through self-supervised learning.
During the preparation of this work, the author(s) used ChatGPT and Quillbot in order to ensure grammatical correctness of the text. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.

### v2

@@
 Abstract
-A common dilemma while photographing a scene is whether to capture it in wider angle, allowing more of the scene to be covered but in lesser details or to click in narrow angle that captures better details but leaves out portions of the scene.
+A common dilemma while photographing a scene is whether to capture it at a wider angle, allowing more of the scene to be covered but in less detail or to click in a narrow angle that captures better details but leaves out portions of the scene.
 We propose a novel method in this paper that infuses wider shots with finer quality details that is usually associated with an image captured by the primary lens by capturing the same scene using both narrow and wide field of view (FoV) lenses.
-We do so by training a GAN-based model to learn to extract the visual quality parameters from a narrow angle shot and to transfer these to the corresponding wide-angle image of the scene.
+We do so by training a Generative Adversarial Network (GAN)-based model to learn to extract the visual quality parameters from a narrow-angle shot and to transfer these to the corresponding wide-angle image of the scene using residual connections and an attention-based fusion module.
 We have mentioned in details the proposed technique to isolate the visual essence of an image and to transfer it into another image.
@@
 keywords:
-Ultra-wide Image , Super Resolution , Generative Adversarial Network
-[…]
-[auth1]organization=Department of Computer Science and Engineering, University of Dhaka, city=Dhaka
-[…]
-[auth2]organization=Department of Computer Science and Engineering, Jahangirnagar University, city=Dhaka
-[…]
-[auth3]organization=Department of Computer Science and Engineering, University of Dhaka, city=Dhaka
+Ultra-wide Image, Super Resolution, Generative Adversarial Network, Attention-based Fusion
 1 Introduction
-Ultra-wide photography is a popular photographic trend in today’s world.
-This form of image-taking, though accommodating a greater field of view and thus offering a unique aesthetic, degrades the finer details better retained by the primary lens.
-This trade-off between the field of view covered and the quality of the finer details retained by the captured image places the users in an unwanted dilemma where one has to compromise the field of view for the details or vice-versa.
-Our primary goal in this paper is to enable the camera users to capture wide-angle images with quality that is identical to the narrow-angle shots while covering the original wide field of view.
-The fundamental idea of capturing wide-angle images that retain the finer details associated with narrow-angle shots is to capture the scene of interest with two different lenses, both narrow and wide.
-The expectation here is that the narrow-angle shot provides the expected resolution while the wide-angle one carries the contents of the scene not covered in the narrow-angle counterpart.
-The idea would then be to render the contents captured by the wide-angle lens in the resolution that is represented by the narrow-angle counterpart.
-As it has been shown by the paper on Neural Style Transfer [6], the style and content of an image are separable and the stylistic representation of an image can be transferred into another image.
-We have leveraged this finding to transfer the resolution qualities of the narrow-angle shot into the wide-angle counterpart.
-Furthermore, we have implemented a pipeline that incrementally transfers the qualitative representations from the narrow-angle shot to the immediate wide-angle shot and repeats the process until the widest available image (often 0.5x or ultra-wide image) is reached.
-This involves capturing the scene of interest using all available lenses and incrementally applying the finest available quality (from the image with the narrowest field of view) through the stack of available images up to the image with the widest field of view.
-In this paper, we propose a GAN-based architecture where the generator network produces realistic-looking enhanced versions of the wide FoV image while the discriminator component attempts to identify generated high-resolution results from the ground truth.
-In the first step, the spatial resolution of the original wide FoV image is upscaled using bicubic interpolation so that more details can be physically accommodated.
-Once upscaled, both the enhanced wide (from now onwards to be referred to as the wide FoV) and the narrow FoV images are partitioned into smaller patches so that visually similar patches may be utilized to assist the process of resolution transfer, inspired by the paper on Vision Transformer [11].
-For the purpose of resolution transfer, we propose a modified version of the Gram matrix representation of an image, as used in the paper on Neural Style Transfer.
-Upon transferring the resolution from the narrow FoV patches into that from the wide FoV image, the synthetic image is passed through the discriminator setup that refines the seams between the tiles and the overall perceptual fidelity of the image, leveraging the task-specific loss functions we have designed.
-Finally, we perform this routine of resolution transfer throughout the stack of all available lenses to attain the highest possible resolution for the widest available field of view.
-Our contributions can thus be summarized into the following points,
-  • 1.
-    An end-to-end pipeline capable of transferring finer quality details to the wide-angle image from the corresponding narrow-angle image
-  • 2.
-    The novel concept of transferring visual quality across the stack of all available lenses
-  • 3.
-    An innovative modification of the Gram matrix that is able to capture the visual essence of an image
+Modern smartphones and mirrorless cameras increasingly incorporate ultra-wide lenses [1] to capture expansive scenes in a single frame.
+While these lenses deliver a striking field of view, they often struggle to preserve the fine-grained detail achieved by conventional “primary” lenses, as discussed in the paper [2].
+Users face a perennial trade-off: opting for a wider perspective at the expense of sharpness or retaining detail by narrowing the view.
+In this work, we seek to resolve this dilemma by synthesizing wide-angle images whose perceptual quality matches that of narrow-angle shots, without sacrificing the original field of view.
+Our approach hinges on jointly leveraging both narrow- and wide-angle captures of the same scene.
+The narrow-angle image supplies high-resolution texture and detail, while the wide-angle image carries context beyond the primary lens’s range.
+We introduce an end-to-end pipeline that incrementally transfers visual fidelity from the finest (narrowest) view through to the widest view in a multi-lens stack.
+First, we partition both the wide image and the corresponding narrow-angle image into spatially aligned patches inspired by the paper on Vision Transformer [3].
+Each patch pair is fed into a generator, which fuses content and detail via cross-view attention and further super-resolves the wide patch using pixel-shuffle layers—progressively enhancing resolution and texture across the full frame.
+Central to our method is a novel GAN [4] architecture: a generator tailored to fuse cross-view content and detail features through attention mechanisms and a discriminator designed to enforce both global realism and seamless patch blending.
+We further extend the classic Gram matrix representation to encode not only texture but also color, brightness, contrast, and sharpness metrics, ensuring rich visual fidelity.
+Applying this procedure iteratively across all available zoom levels—from telephoto to ultra-wide—we end up with a single image that combines the sweeping perspective of the widest lens with the crisp detail of the primary lens.
+Our key contributions are
+  • An end-to-end pipeline that transfers high-resolution detail from narrow to wide-angle images.
+  • A GAN framework with cross-view attention and an enhanced Gram matrix formulation for robust detail propagation.
+  • A cascading lens-stack strategy that scales seamlessly from telephoto to ultra-wide captures.
 2 Related Works
@@
 Later, generative adversarial models such as SRGAN [6], ESRGAN [7], and Real-ESRGAN [8] improved perceptual realism by emphasizing texture fidelity and producing visually convincing outputs.
-However, these models still struggle with preserving structural consistency, especially in scenes where the narrow and wide FOV images exhibit differences in perspective or content.
+However, these models still struggle with preserving structural consistency.
 Another stream of relevant research comes from neural style transfer.
@@
 However, transferring quality between narrow and wide FOV images introduces unique challenges not fully addressed by style transfer methods, as the wide FOV often contains additional contextual elements that require accurate integration.
+Figure 1: Overall model architecture
 Recent advancements have introduced lightweight and efficient architectures tailored for real-time applications.
@@
 This design facilitates the reconstruction of high-quality images without the overhead associated with deeper networks.
-Building upon the success of attention mechanisms, the Efficient Long-Range Attention Network (ELAN) [25] integrates shift convolutions with group-wise multi-scale self-attention modules.
+Building upon the success of attention mechanisms, the Efficient Long-Range Attention Network (ELAN) [13] integrates shift convolutions with group-wise multiscale self-attention modules.
 This combination captures both local and long-range dependencies, enhancing the model’s ability to reconstruct fine details across varying scales.
-ELAN’s architecture achieves a balance between performance and efficiency, making it suitable for applications requiring real-time processing.
+The ELAN architecture achieves a balance between performance and efficiency, making it suitable for applications that require real-time processing.
 Further exploring attention-based models, DLGSANet [14] introduces a hybrid approach combining dynamic local and sparse global self-attention mechanisms.
@@
 This dual attention strategy enhances the model’s capacity to reconstruct high-resolution images with reduced computational demands.
-In parallel, the Deep Multi-level Feature Fusion Network (DMFFN) [26] addresses common challenges in super-resolution tasks, such as blurred edges and redundant structures.
-By employing two-level nested residual blocks with asymmetric structures, DMFFN efficiently fuses multi-level features, leading to improved edge preservation and faster convergence during training.
-These contemporary models demonstrate significant progress in balancing performance and efficiency, particularly in scenarios requiring the enhancement of wide FOV images by leveraging information from narrow FOV counterparts.
+In parallel, the Deep Multilevel Feature Fusion Network (DMFFN) [15] addresses common challenges in super-resolution tasks, such as blurred edges and redundant structures.
+Using nested two-level residual blocks with asymmetric structures, DMFFN efficiently fuses multilevel features, leading to improved edge preservation and faster convergence during training.
+These contemporary models demonstrate significant progress in balancing performance and efficiency.
 Their architectures are designed to maintain structural consistency and fine details, addressing the limitations observed in earlier methods.
-Moreover, the integration of perceptual metrics such as LPIPS [16] has become increasingly prevalent, offering a more nuanced assessment of visual similarity that aligns closely with human perception.
-Incorporating human-in-the-loop evaluations further ensures that the enhanced images meet real-world expectations for quality and accuracy.
-In this work, building upon these advancements, we aim to improve wide FOV image quality by transferring visual fidelity from narrow FOV versions of the same scene.
+Perceptual metrics such as LPIPS [16] have become widely adopted for evaluating visual similarity in a way that aligns closely with human perception.
+Beyond perceptual loss, incorporating visual cues that guide perception—namely color, brightness, contrast, and sharpness—is equally important.
+We select these attributes based on extensive literature: [17] identifies them as key appearance properties in image quality assessment; [18] highlights the role of color in quality estimation; [19] emphasizes the critical balance of brightness and contrast in enhancement; and [20] combines contrast and sharpness to assess enhancement reliability.
 By combining insights from generative models, reference-based super-resolution, attention mechanisms, and perceptual consistency techniques, we target high-quality, structurally accurate image synthesis suited for immersive visual applications.
 3 Method
-In this section, we have described in detail the adopted methodologies and architecture of the model used to transfer high-quality details from the narrow-angle image into the wide-angle counterpart.
-Figure 1: Model Architecture
-In the case of training the model, we consider the original image as the wide FoV ground truth and extract the central portion to serve as the high-definition narrow FoV.
-The original wide FoV is then downsampled to simulate the captured wide FoV image.
-3.1 Content Reconstruction
-Our primary goal is to reconstruct the contents of the wide-angle image applying the visual qualities of the narrow-angle counterpart.
-By visual qualities we indicate the parameters that impact the perceptual cues which are color, brightness, contrast, and sharpness.
-For reconstructing the contents of the wide-angle image, we utilize the pixel-wise content loss.
-We predict the pixel entries for a white-noise image such that the following content loss is minimized.
-$L^{(l)}_{content}(p,q)=\frac{1}{2}\sum_{i,j}(Q^{l}_{ij}-P^{l}_{ij})^{2}$
-Here $p$ and $q$ are the original and white-noise images respectively and $P^{l}_{ij}$ and $Q^{l}_{ij}$ are the feature representations in the $l^{th}$ layer for $i^{th}$ filter at the position $j$.
-3.2 Visual Quality Reconstruction
-As mentioned earlier, the visual qualities are obtained from the narrow-angle image.
-The qualities are extracted from the image using the gram matrix approach.
-We have divided the image into a fixed number of patches and computed Gram matrices for each patch.
-The motivation to adopt this approach comes from the intuition that visually similar patches may exist between the narrow and the wide-angle images.
-Obtaining visual qualities from the patch of narrow-angle image that visually aligns with a patch of the wide-angle image would produce better results.
-Hence, we divide the entire narrow and wide-angle images into a predefined number of patches.
-For each patch of the wide-angle image, the patch from the narrow-angle image that is visually most similar is selected.
-The visual similarity score between patches is determined using a perceptual loss function described in the following section.
-Patches are taken to be visually similar if the similarity score exceeds a threshold.
-In case no patch from the narrow-angle image is found to exceed the threshold, a gram matrix representation of the entire narrow-angle image is used as a backup.
-Figure 2 illustrates partitioning the image into fixed-sized patches.
-Images partitioned into fixed-sized patches just for illustrative purposes do not indicate the actual number of patches or their proportions used in the implementation.
-The patches are bound by black lines.
-Arrows and bounding boxes of different colors denote patches in the narrow FoV image that are visually similar to certain patches from the wide FoV image.
+The core philosophy of our solution is to use a cross-view attention mechanism to fuse content and visual quality features from a given pair of wide and narrow-angle images.
+The images are initially partitioned into fixed-size patches.
+We pick a pair of wide and narrow patches from the pool and pass them across the network architecture, as described below.
+3.1 Image Partitioning
+Both narrow and wide FoV images are partitioned into $128\times 128$-sized patches to ensure portions of images sharing visual similarity are used to exchange content and visual quality data.
+Once partitioned, patches from wide and narrow FoV images that share visual properties are mapped using perceptual loss [21] and stored in the pool, ready to be pushed into the generator network.
+A cosine similarity on the feature maps generated by the mid layers of a VGG-19 [22] backbone is used to determine the perceptual distance between patches.
 Figure 2: Image partitioning and patch similarity between narrow and wide FoV images.
-3.2.1 Dividing into Patches
-The narrow and wide-angle images are first divided into patches of dimension $128\times 128$.
-This particular size for the patches has been chosen to create a balance between structural information contained within an image and the computational cost incurred to process the patches.
-The idea of dividing the images into patches also takes inspiration from the paper on Vision Transformer [11].
-3.2.2 Patch-wise Similarity
-We use perceptual loss [17] to quantify the visual similarity between patches from narrow and wide-angle images.
-We use the mid layers of a VGG19 network to numerically determine the perceptual loss between the pair of patches.
-Following function is used to determine the loss,
-$L_{feature}(p_{1},p_{2})=\frac{1}{C_{j}*H_{j}*W_{j}}||\phi_{j}(p_{2})-\phi_{j}(p_{1})||_{2}^{2}$
-Here $p_{1}$ and $p_{2}$ are the patches from wide and narrow-angle images, $\phi_{j}(p_{1})$ and $\phi_{j}(p_{2})$ denote the feature map obtained from the $j^{th}$ layer, and $C_{j},H_{j}$ and $W_{j}$ are the number of channels, height, and width of the patches.
-3.3 Generator Architecture
-Figure 3: Generator Architecture
-We utilize a ResNet [27] based generator with residual blocks.
-The input layer is followed by a collection of residual blocks to facilitate earlier layers in learning complex functions.
-The residual block relays the data into the upsampling block that generates the enhanced patch and passes through a final convolution layer to adjust the channel dimension.
-3.3.1 Feature Map Construction
-We consider the following parameters to effectively represent the visual qualities of an image (patch of an image in our case); $color,contrast,brightness,sharpness$.
-Considering each of these parameters as distinct features of a patch, an enhanced feature map representation of the patch is computed.
-The enhanced feature map is an augmentation of the individual features extracted from the patch and is formulated as follows,
-$\hat{F_{l}}=[F_{l},F_{l_{color}},F_{l_{brightness}},F_{l_{contrast}},F_{l_{sharpness}}]$
-Where $F_{l}\epsilon\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ is the feature map extracted using a VGG backbone network. Rest of the features are computed as follows,
+Figure 3: Detailed architecture for the Generator component
+3.2 Network Architecture
+The architecture of the generator network is built upon a stack of ResNet-like [23] residual blocks that take as inputs a pair of wide and narrow FoV patches from the pool.
+The patches are encoded using separate encoder components to capture structural (content) and visual features from wide and narrow FoV patches respectively.
+Each block then employs a cross-view attention mechanism to fuse the content and visual properties from wide and narrow feature maps respectively.
+The pipeline consists of the following major components,
+3.2.1 Structural Feature Encoder
+Wide FoV patches are encoded with traditional CNN-based encoders to capture the content details. Encoded wide patch represented as $E_{w}$
+3.2.2 Visual Feature Encoder
+These encoders are specialized in extracting the visual parameters from a patch and are used to encode the narrow FoV patches.
+We use the Gram matrix approach to represent the visual features of a patch and include the following additional parameters in the Gram matrix representation: color, brightness, contrast, and sharpness.
+The feature map is constructed as follows,
+  • The original feature map $F_{l}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ is generated using a VGG-19 network.
+  • Color information for each channel over spatial dimensions, also referred to as global average pooling [24], is computed as:
+    $[F_{l_{\text{color}}}]_{c}$ $=\frac{1}{H_{l}W_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}F_{l}(c,h,w),$
+    for $\quad\text{for }c=1,\dots,C_{l}$
+  • Global mean pixel intensity across all channels and spatial dimensions indicates the perceived luminance (brightness) of the patch. Computed as:
+    $F_{l_{\text{brightness}}}=\frac{1}{C_{l}H_{l}W_{l}}\sum_{c=1}^{C_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}F_{l}(c,h,w)$
+  • We represent the overall contrast of the patch using the standard deviation:
+    $F_{l_{\text{contrast}}}=\sqrt{\frac{1}{C_{l}H_{l}W_{l}}\sum_{c=1}^{C_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}\left(F_{l}(c,h,w)-\mu_{l}\right)^{2}}$
+    where $\mu_{l}=F_{l_{\text{brightness}}}$ is the global mean defined above.
+  • Sharpness is accounted for using the Laplacian operated, as mentioned in this paper [25]:
+    $F_{l_{\text{sharpness}}}(c,h,w)=(F_{l}(c,\cdot,\cdot)*\nabla^{2})(h,w)$
+    where * denotes the convolution and $\nabla^{2}$ is the Laplacian kernel.
+  • Finally, we augment the original feature map obtained from VGG-19 as follows:
+    $\hat{F}_{l}=[F_{l},F_{l_{\text{color}}},F_{l_{\text{brightness}}},F_{l_{\text{contrast}}},F_{l_{\text{sharpness}}}]$
+    where scalar or vector features are broadcast to match the spatial dimensions of $F_{l}$.
+Once the enhanced feature map has been computed, we perform a dot product between the enhanced feature map and its transpose to attain the Gram matrix representation $E_{n}$, carrying the visual essence of the narrow FoV patch.
+$E_{n}=\hat{F}_{l}\hat{F}_{l}^{\top}$
+3.2.3 Cross-Attention Residual Blocks
+Once the patches are encoded into the required latent space, we employ an attention-based fusion mechanism that fetches the most relevant visual features from the narrow patch and fuses those with the content features generated from the wide patch.
+  • Encoded wide FoV patch $E_{w}$, representing the content data, acts as the query matrix
+  • Encoded narrow FoV patch $E_{n}$, carrying the visual properties, provides both the key and value matrices
+  • The attention scores, which gather semantically aligned, superior quality visual features present in the encoded narrow FoV patch, are computed using the scaled dot product equation introduced in the attention paper [26],
+    $Attention(E_{n},E_{w})=softmax(\frac{E_{w}E_{n}^{\top}}{\sqrt{d_{k}}})E_{n}$
+  • The gathered attention scores are then incorporated into the wide FoV feature map using a residual addition, producing the feature map representation of the enhanced wide FoV patch, $E$
+    $E=E_{w}+Attention(E_{n},E_{w})$
+3.2.4 Upsampling Layers
+The output of the final residual block is passed through a pixel shuffle [27] layer with a scale factor equal to 2 to super-resolve the patch, denoted by $E^{*}$.
+3.3 Loss Calculation
+Generator loss is calculated patch-wise, i.e., each of the generated patches is compared against both the original wide and narrow FoV patches to quantify the content and visual deviations from the ground truth.
+Then the patches are assembled using seamless cloning, to eliminate visible seams during blending, and the reconstructed image is passed through the discriminator which computes a seam consistency loss over the entire image.
+The losses are described as follows,
+3.3.1 Generator Loss
   1. 1.
-    Color: An average of the RGB channels forms an embedding that represents the overall intensities. Computed as,
-    $F_{l_{color}}=\frac{1}{H_{l}W_{l}}\sum^{H_{l}}_{i=1}\sum^{W_{l}}_{j=1}F_{l}$
+    Content Loss: We find the deviations of the contents of the generated patch and ground truth with the following pixel-wise difference,
+    $L^{(l)}_{content}(E^{o}_{w},E^{g}_{w})=\frac{1}{n}\sum_{i}(E^{o}_{w}-E^{g}_{w})^{2}$
+    Here $E^{o}_{w}$ and $E^{g}_{w}$ are the latent space representations of the original and generated patches.
   2. 2.
-    Brightness: Mean pixel intensities over all pixels are taken to indicate the perceived luminance of the patch. Computed as,
-    $F_{l_{brightness}}=\frac{1}{C_{l}H_{l}W_{l}}\sum^{C_{l}}_{i=1}\sum^{H_{l}}_{i=1}\sum^{W_{l}}_{k=1}F_{l}$
+    Visual Quality Loss: This loss is calculated as the difference of the Gram matrix representation of the generated patch and the narrow FoV patch using the following equation,
+    $L^{(l)}_{visual}=\sum_{l}w_{l}||E^{o}_{n}-E^{g}_{n}||^{2}$
+    Here $E^{g}_{n}$ and $E^{o}_{n}$ are the gram matrix representations of the patches from the generated and the original narrow FoV patches, respectively.
   3. 3.
-    Contrast: Variance of the pixel intensities is computed to represent the variation of intensities, that is the contrast, within the patch. Calculated as,
-    $F_{l_{contrast}}=\sqrt{\frac{1}{C_{l}H_{l}W_{l}}\sum^{C_{l}}_{i=1}\sum^{H_{l}}_{i=1}\sum^{W_{l}}_{k=1}(F_{l}-\mu_{l})^{2}}$
-    here $\mu_{l}$ is the average intensity.
-  4. 4.
-    Sharpness: Sharpness is captured using the Laplacian filter. The Laplacian filter is multiplied to the feature map as follows,
-    $F_{l_{sharpness}}=F_{l}*\bigtriangledown^{2}$
-3.3.2 Gram Matrix Computation
-Once the feature map is constructed, we can use this representation to find the gram matrix representation of the patch of the image for the current layer. The computation can be performed as follows,
-$G_{l}=\hat{F}_{l}\hat{F}_{l}^{\top}$
-The gram matrix is the product of the enhanced feature map to its transpose, essentially capturing the pair-wise correlation between features.
-3.3.3 Loss Calculation
-Visual quality loss is defined as the squared distance between the gram matrix representation of the source and target images.
-In our case, the source is the patch from the narrow FoV image while the target is a white-noise image which is eventually going to ”be” the resultant image by minimizing the loss defined above.
-The numerical representation of the loss is,
-$L^{(l)}_{visual}=\sum_{l}w_{l}||G_{l_{q}}-G_{l_{p}}||^{2}$
-Here $G_{l_{q}}$ and $G_{l_{p}}$ are the gram matrix representations of the patches from the white-noise and the narrow FoV images respectively.
-Mentionable that in case no particular match is found for a certain wide FoV patch in the narrow FoV, $G_{l_{p}}$ would denote the gram matrix representation of the entire narrow FoV image.
-Hence the total generator loss is quantified as,
+    Total generator loss can be written as,
     $L^{(l)}_{G}=L^{(l)}_{content}+L^{(l)}_{visual}$
-3.4 Discriminator Loss
+3.3.2 Discriminator Loss
 For the generated output image to be realistic and visually plausible, the results generated by the deconvolution network is passed through a discriminator component that aims to differentiate synthetic images from the real ones.
@@
 The losses are described as follows,
-3.4.1 Seam Consistency Loss
-This loss is aimed at eliminating the seams between the patches in the synthetic image by minimizing the perceptual difference at the boundaries of the adjacent patches.
+  1. 1.
+    Seam Consistency Loss This loss is aimed at eliminating the seams between the patches in the synthetic image by minimizing the perceptual difference at the boundaries of the adjacent patches.
 For adjacent patches $P_{i}$ and $P_{j}$, we take neighboring regions $R_{i}$ and $R_{j}$ which are $w$ pixels wide and compute the loss as,
     $L^{(l)}_{seam}=\frac{1}{w}\sum||\phi_{l}(R_{j})-\phi_{l}(R_{i})||^{2}$
-3.4.2 Perceptual Loss against Ground Truth
-To ensure the essence of the narrow FoV has been plausibly transferred to the synthetic image, the output is perceptually compared to the original wide FoV ground truth.
+  2. 2.
+    Perceptual Loss against Ground Truth To ensure the essence of the narrow FoV has been plausibly transferred to the synthetic image, the output is perceptually compared to the original wide FoV ground truth.
 We use the same VGG19 activation maps to find the perceptual difference $L^{(l)}_{perceptual}$ between the generated wide FoV image and the original image from the dataset.
@@
 4.1 Datasets
-We have used two datasets to ensure robust training of our proposed model.
-  • 1.
-    DIV8K [19]: High-resolution images widely used for super-resolution tasks.
-More than 1500 photos with a maximum resolution of 8K are included in the dataset.
-It covers a wide range of scene contents.
-  • 2.
-    Landscapes Dataset [18] (LHQ 1024×1024): With 90,000 high-quality images of 1024 x 1024 resolution, it includes a diverse range of urban and natural environments.
+We have used Landscapes Dataset [28] (LHQ 1024×1024) to train the proposed model.
+With 90,000 high-quality images of 1024 x 1024 resolution, it includes a diverse range of urban and natural environments.
 These high-quality landscape images with rich textures and details challenge the method’s ability to handle complex natural scenes.
-4.2 Generator
-As stated earlier, the generator network takes in $128\times 128\times 3$ patches and performs an initial convolution followed by a ReLU layer.
-The feature map is then fed into an array of $64$ residual blocks.
-Residual blocks assist in preventing the gradient from dying out during back propagation.
-This is followed by $2$ upsampling blocks, responsible for physically upscaling the patch using Pixel Shuffle layers.
-The upsampled patches passed through the output convolution block to adjust the channel dimension and come out as enhanced patches with a dimension of $256\times 256\times 3$.
-4.3 Discriminator
-We use a ResNet18 based discriminator network that takes as input a $1024\times 1024\times 3$ image formed by stitching the enhanced patches generated by the generator using Poisson Blending.
-The image undergoes a series of $6$ residual blocks followed by a layer of adaptive pooling and a fully connected layer outputting a single scaler denoting the realness of the input image.
+We have randomly sampled 20000 + 500 images (discussed in the Training Details section) and fed into the training loop in the ascending order of pixel variance [29], to ensure “easier” images are encountered first.
+Figure 4: Convergence of loss curves shown over multiple iterations
+[H] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 2x Zoom 1x Zoom 1x Output [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 1x Output 0.5x Zoom 0.5x Output
+Table 1: Experimental results demonstrating the cascading effect of multi-zoom image upscaling.
 4.2 Simulating Narrow and Wide FoVs
-A high-quality narrow FoV and two wide FoV (one low-quality and one high-quality) images are required for training.
-High-quality wide images are used to calculate the loss, whereas low-quality wide and high-quality narrow images are used as model inputs.
-The original images from the dataset are treated as high-quality wide FoV images as well as the ground truth.
-To simulate high-quality narrow and low-quality wide-angle images from single-image datasets we have used the following techniques.
 High-quality narrow FoV simulation: For each image in the dataset, we extract a high-resolution central region with a zoom factor of $\frac{5}{3}$.
@@
 The resulting image replicates the characteristics of a narrow FoV while maintaining the high quality of the original.
-Low-quality wide FoV simulation: For the DIV8K dataset with 8K resolution, we simulated low-quality wide FoV images by first downsampling each image to 4096 × 4096 and then upsampling it back to 8K.
-Similarly, for the Landscapes dataset (LHQ 1024 × 1024), we downsampled each image to 512 × 512 and then upsampled it back to 1024 × 1024.
-This intentional degradation process mimics low-quality imaging conditions while preserving the broader scene context, allowing us to test our methods under challenging visual conditions.
-The description of workflow provided in this paper is in light of the Landscape dataset, that is, assuming the original image to be 1024 x 1024.
-The image reconstruction pipeline has been shown in the following Figure-4.
+Low-quality wide FoV simulation: We downsampled each image by applying a blur kernel and adding noise following this paper [30], as well as resizing to 512 × 512, and then upsampled it (using bicubic interpolation) back to 1024 × 1024.
+This intentional degradation process mimics the wide image of low-quality imaging conditions while preserving the broader scene context, allowing us to test our methods under challenging visual conditions.
+4.3 Generator
+The generator takes in a pair of patches, each of dimension $128\times 128\times 3$.
+The wide FoV patch is passed through a CNN encoder with a final batchNorm layer to produce the content feature map.
+The narrow FoV patch initially undergoes feature extraction through the $Conv2_{2},Conv3_{2},Conv4_{2}$ layers of a VGG-19 backbone.
+The feature map produced by VGG-19 is enhanced by the visual properties computed from the original patch.
+We use 10 attention-based residual blocks, each with two $3\times 3$ convolution units, a ReLU, followed by the cross-view attention score calculation unit, outputs of which are added to the wide FoV patch.
+The details-infused wide FoV patch is then passed through a series of convolution layers interleaved by pixel shuffle layers to physically upscale the patch to a dimension of $256\times 256\times 3$.
+4.4 Discriminator
+We use a ResNet18 [31] based discriminator network that takes as input a $1024\times 1024\times 3$ image formed by stitching the enhanced patches generated by the generator using Poisson Blending [32].
+The image undergoes a series of $6$ residual blocks followed by a layer of adaptive pooling and a fully connected layer outputting a single scalar denoting the realness of the input image.
+4.5 Enhanced image reconstruction pipeline
+Figure 5 highlights the steps undertaken to reconstruct the enhanced.
 Figure 5: Enhanced image reconstruction pipeline
@@
 Key hyperparameters for training and evaluation included a patch size of 128 × 128, Adam optimizer with a learning rate of 0.0001 with exponential decay every 10 epochs, a batch size of 16, perceptual loss layers (Conv2_2, Conv3_2, Conv4_2 of VGG19), and a threshold for patch similarity empirically set to 0.7.
-The generator network was pre-trained for 50 epochs using content and visual losses before incorporating adversarial loss.
-The full generator-discriminator framework was then trained for an additional 200 epochs, totaling 400K iterations.
-The generator consists of 16 residual blocks, each with two 3×3 convolutional layers, batch normalization, and parametric ReLU activation, allowing effective feature extraction while keeping training stable.
-To enhance high-frequency details, two upsampling layers using pixel shuffle progressively increase the resolution.
-The final 3×3 convolutional layer reconstructs the high-resolution image.
-The convergence behavior of the visual quality loss for a single patch in the generator network is shown in Figure 5.
-It highlights the effectiveness of our proposed enhanced Gram matrix, which causes an earlier convergence by approximately 30K iterations compared to the traditional approach.
-This earlier stabilization suggests that our method improves the efficiency of optimization, reducing the training time required while maintaining reconstruction quality.
-Figure 5: Faster convergence of visual quality loss with the enhanced Gram matrix.
-To further stabilize the training and prevent mode collapse, spectral normalization was applied to the discriminator while gradient clipping controlled exploding gradients.
-The model was optimized using a combination of content loss (L1), perceptual loss (VGG-based), and adversarial loss, ensuring a strong balance between detail accuracy and perceptual quality.
-4.6 Experiment Results
-Table 1 presents the results of our experiment, where a narrow and a low-resolution wide image are used as input. These images are processed through our model to generate the final output.
-Table 2 illustrates the cascading effect of the outputs derived from images captured at different zoom levels.
-In the first row, the input images consist of a 2x (telephoto) and a 1x (standard) zoom, which collectively produce a high-resolution 1x (standard) output.
-The second row takes the output from the previous row as one input and combines it with a low-resolution 0.5x (wide) image, ultimately generating a high-resolution 0.5x (wide) image.
-This sequence effectively simulates how our method can be applied in a cascading manner using all available zoom lenses.
+The generator network was pre-trained for 50 epochs using content and visual losses before incorporating adversarial loss, each epoch carrying 10 sample images.
+The full generator-discriminator framework was then trained for an additional 2000 epochs, totaling 20.5K iterations.
+The convergence behavior of the generator loss has been illustrated in Figure 4.
+These curves represent the first 15 epochs where the generator has been trained solo.
+Upper row features the raw loss curves as generated by the GAN architecture, while the lower row has been smoothened using the numpy convolve [33] method.
 Narrow FoV Input wide FoV Our result
@@
 Table 2: Experimental results showing the output generated by our model when processing narrow and wide input images.
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-2x Zoom 1x Zoom 1x Output
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-1x Output 0.5x Zoom 0.5x Output
-Table 2: Experimental results demonstrating the cascading effect of multi-zoom image upscaling.
-5 Comparison with Prior Arts
-We have evaluated our proposed methods on PSNR, SSIM and LPIPS[16] metrics over Set5[12], Set14[13] and BSD100[14] datasets.
-For PSNR and SSIM, we have compared our results with existing technologies such as ESRGAN[2], LatticeNet[24], ELAN[25], DLGSANet[23] and DMFFN[26] in Table 3 and Table 4.
+[H] Dataset bicubic Multiscale GAN Ours Set5 0.5657 0.2823 0.0795 Set14 0.6114 0.3717 0.2032 BSD100 0.5726 0.4112 0.3361
+Table 3: LPIPS Comparison (best marked in red)
+Figure 6: Visual comparison of our results against the mentioned prior arts
 Dataset ESRGAN LatticeNet ELAN DLGSANet DMFFN Ours
@@
 Table 5: SSIM Comparison (highest marked in red)
+4.7 Experiment Results
+Table 2 presents the results of our experiment, where a narrow and a low-resolution wide image are used as input. These images are processed through our model to generate the final output.
+Table 1 illustrates the cascading effect of the outputs derived from images captured at different zoom levels.
+In the first row, the input images consist of a 2x (telephoto) and a 1x (standard) zoom, which collectively produce a high-resolution 1x (standard) output.
+The second row takes the output from the previous row as one input and combines it with a low-resolution 0.5x (wide) image, ultimately generating a high-resolution 0.5x (wide) image.
+This sequence effectively simulates how our method can be applied in a cascading manner using all available zoom lenses.
+5 Comparison with Prior Arts
+We have evaluated our proposed methods on PSNR, SSIM and LPIPS [16] metrics over Set5 [34], Set14 [35] and BSD100 [36] datasets.
+For PSNR and SSIM, we have compared our results with existing technologies such as ESRGAN [7], LatticeNet [12], ELAN [13], DLGSANet [14] and DMFFN [15] in Table 4 and Table 5.
 We have presented our comparison with bicubic interpolation and Multiscale GAN [37] on the LPIPS metric in Table 3.
-It is to be noted that LPIPS values are within the range $[0,1]$, with a lower value denoting the generated image is perceptually more aligned with the ground truth.
-Dataset bicubic Multiscale GAN Ours
-Set5 0.5657 0.2823 0.0795
-Set14 0.6114 0.3717 0.2032
-BSD100 0.5726 0.4112 0.3361
-Table 5: LPIPS Comparison (best marked in red)
-Figure 6 presents a comparative analysis of our method against several prior super-resolution models, including bicubic interpolation, LatticeNet [24], ELAN [25], and DLGSANet [23].
-The leftmost image represents the original ground truth, with the red bounding box highlighting the region of interest for focused evaluation.
-The comparative results illustrate variations in texture reconstruction, detail preservation, and structural fidelity across different approaches.
-Notably, our method, displayed in the bottom-right corner, achieves superior sharpness and consistency, effectively reducing blurring and artifacts observed in other models.
-Figure 6: Visual comparison of super-resolution models, including bicubic interpolation, LatticeNet, ELAN, and DLGSANet.
-The red bounding box highlights the region of interest, with our method (bottom-right) achieving superior detail preservation and structural fidelity
+Visual comparison of our method against prior super-resolution models has been presented in Figure 6.
 6 Conclusion
-We proposed a GAN-based approach to enhance wide-angle images by transferring fine details from a narrow-angle shot.
-Our method balances coverage and detail through multi-zoom image fusion, leveraging perceptual losses and an improved Gram matrix representation.
-The cascading lens stack further extends its application to multi-lens imaging systems.
-Extensive experiments showed that our approach outperforms existing super-resolution and image synthesis methods, achieving superior PSNR, SSIM, and LPIPS scores while reducing artifacts.
-This opens new possibilities for computational photography, smartphone imaging, and digital zoom.
-Future work will focus on improving performance under extreme lighting, dynamic scenes, and real-time constraints, with potential enhancements through self-supervised learning.
-During the preparation of this work, the author(s) used ChatGPT and Quillbot in order to ensure grammatical correctness of the text.
-After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.
+In this work, we presented a novel approach for enhancing ultra-wide field-of-view (FoV) images by leveraging paired narrow-FoV views of the same scene.
+The core of our method is a patch-based cross-view attention mechanism that selectively transfers fine-grained details from narrow-FoV patches to their wide-FoV counterparts, guided by Gram matrix-based visual encoding that captures texture statistics for accurate alignment.
+To further refine the reconstruction, we introduce a cascading lens stack mechanism that progressively transfers details across adjacent lens images, facilitating spatial consistency and coherent integration across the full image.
+We trained our model on the Landscape HQ dataset to ensure coverage of both structural diversity and fine detail.
+For evaluation, we used standard benchmark datasets including BSD100, Set5, and Set14.
+Experimental results demonstrate that our method consistently outperforms strong baselines such as ESRGAN, ELAN, DLGSANet, DMFFN, and LatticeNet in both PSNR and SSIM metrics.
+Qualitative comparisons further confirm the perceptual superiority of our results.
+Overall, our framework successfully addresses the challenge of enhancing ultra-wide FoV images by combining cross-view attention, texture-aware encoding, and a cascading lens-based reconstruction strategy.
+Future extensions could explore its application to video enhancement and real-time inference scenarios.

### v3

@@
-Enhancing Wide-Angle Image Using Narrow-Angle View of the Same Scene
+
 Abstract
-A common dilemma while photographing a scene is whether to capture it at a wider angle, allowing more of the scene to be covered but in less detail or to click in a narrow angle that captures better details but leaves out portions of the scene.
-We propose a novel method in this paper that infuses wider shots with finer quality details that is usually associated with an image captured by the primary lens by capturing the same scene using both narrow and wide field of view (FoV) lenses.
-We do so by training a Generative Adversarial Network (GAN)-based model to learn to extract the visual quality parameters from a narrow-angle shot and to transfer these to the corresponding wide-angle image of the scene using residual connections and an attention-based fusion module.
-We have mentioned in details the proposed technique to isolate the visual essence of an image and to transfer it into another image.
-We have also elaborately discussed our implementation details and have presented the results of evaluation over several benchmark datasets and comparisons with contemporary advancements in the field.
+A common dilemma in photography is the trade-off between capturing a wider field of view with less detail and a narrower field of view with finer detail.
+To address this, we propose a method that enhances wide-angle images by leveraging co-captured narrow-angle views of the same scene.
+Our approach introduces a narrow field-of-view (FoV) image-informed enhancement strategy that overcomes the limitations of blind super-resolution, which often struggles to recover high-frequency details without auxiliary guidance.
+We train a GAN-based model to extract fine-grained visual features from the narrow-view image and fuse them into the corresponding wide-view image through residual connections and an attention-based feature integration module.
+This fusion is guided by perceptual cues representing the visual essence—such as color, contrast, and sharpness—learned from the narrow FoV.
+We describe our architectural design and implementation in detail and demonstrate the effectiveness of our method through extensive experiments and comparisons with recent state-of-the-art techniques.
 keywords:
@@
 Modern smartphones and mirrorless cameras increasingly incorporate ultra-wide lenses [1] to capture expansive scenes in a single frame.
-While these lenses deliver a striking field of view, they often struggle to preserve the fine-grained detail achieved by conventional “primary” lenses, as discussed in the paper [2].
+While these lenses deliver a striking field of view, they often struggle to preserve the fine-grained detail achieved by conventional “primary” lenses.
+As discussed in the paper [2], this difficulty is faced due to various technical issues, such as barrel distortion and spreading the view across a greater number of pixels, resulting in a lower perceived resolution.
 Users face a perennial trade-off: opting for a wider perspective at the expense of sharpness or retaining detail by narrowing the view.
-In this work, we seek to resolve this dilemma by synthesizing wide-angle images whose perceptual quality matches that of narrow-angle shots, without sacrificing the original field of view.
-Our approach hinges on jointly leveraging both narrow- and wide-angle captures of the same scene.
-The narrow-angle image supplies high-resolution texture and detail, while the wide-angle image carries context beyond the primary lens’s range.
-We introduce an end-to-end pipeline that incrementally transfers visual fidelity from the finest (narrowest) view through to the widest view in a multi-lens stack.
-First, we partition both the wide image and the corresponding narrow-angle image into spatially aligned patches inspired by the paper on Vision Transformer [3].
-Each patch pair is fed into a generator, which fuses content and detail via cross-view attention and further super-resolves the wide patch using pixel-shuffle layers—progressively enhancing resolution and texture across the full frame.
-Central to our method is a novel GAN [4] architecture: a generator tailored to fuse cross-view content and detail features through attention mechanisms and a discriminator designed to enforce both global realism and seamless patch blending.
+In this work, we seek to mitigate this dilemma by synthesizing wide-angle images whose perceptual quality matches that of narrow-angle shots, without sacrificing the original field of view.
+While traditional super-resolution (SR) algorithms can increase the spatial resolution of an input image, they must rely on statistical priors and learned patterns to hallucinate high-frequency details, often leading to inconsistent results and visual artifacts.
+Our approach, on the other hand, hinges on jointly leveraging both narrow and wide-angle captures of the same scene.
+The narrow-angle image supplies high-resolution texture and detail, hence supplanting the blind dependence on learned distributions, while the wide-angle image carries context beyond the primary lens’s range.
+The enhancement objective, in this paper, is formulated as a resolution boosting task, informed by fine detail cues from the narrow FoV image, allowing us to draw upon proven strategies and validation procedures from SR literature.
+To that end, we introduce an end-to-end pipeline that utilizes cross-view attention scores to pinpoint the relevant features in the narrow FoV image embedding that can be used to guide the upscaling of the wide FoV counterpart.
+Inspired by the paper on Vision Transformer [3], we adopt a patch-based approach to ensure the quality transfer occurs between regions that share visual similarity.
+Central to our method is a proposed GAN [4] architecture: a generator attempting to fuse cross-view content and detail features through attention mechanisms and a discriminator that aims at enforcing both global realism and seamless patch blending.
 We further extend the classic Gram matrix representation to encode not only texture but also color, brightness, contrast, and sharpness metrics, ensuring rich visual fidelity.
-Applying this procedure iteratively across all available zoom levels—from telephoto to ultra-wide—we end up with a single image that combines the sweeping perspective of the widest lens with the crisp detail of the primary lens.
-Our key contributions are
+Applying this procedure iteratively across all available zoom levels, from telephoto to ultra-wide, our method aims to synthesize the broad perspective of the widest lens while improving detail quality using information from the primary lens.
+Our key contributions are,
   • An end-to-end pipeline that transfers high-resolution detail from narrow to wide-angle images.
@@
 Subsequent work improved efficiency using feed-forward networks [10] and adaptive normalization techniques [11], broadening their use cases.
-However, transferring quality between narrow and wide FOV images introduces unique challenges not fully addressed by style transfer methods, as the wide FOV often contains additional contextual elements that require accurate integration.
+However, transferring quality between narrow and wide FOV images introduces unique challenges not addressed by style transfer methods, as the wide FOV often contains additional contextual elements that require accurate integration.
 Figure 1: Overall model architecture
-Recent advancements have introduced lightweight and efficient architectures tailored for real-time applications.
-LatticeNet [12] employs a lattice block structure inspired by filter banks, enabling adaptive feature fusion while maintaining low computational complexity.
-This design facilitates the reconstruction of high-quality images without the overhead associated with deeper networks.
-Building upon the success of attention mechanisms, the Efficient Long-Range Attention Network (ELAN) [13] integrates shift convolutions with group-wise multiscale self-attention modules.
-This combination captures both local and long-range dependencies, enhancing the model’s ability to reconstruct fine details across varying scales.
-The ELAN architecture achieves a balance between performance and efficiency, making it suitable for applications that require real-time processing.
-Further exploring attention-based models, DLGSANet [14] introduces a hybrid approach combining dynamic local and sparse global self-attention mechanisms.
-The Multi-Head Dynamic Local Self-Attention (MHDLSA) module effectively captures local features, while the Sparse Global Self-Attention (SparseGSA) module focuses on the most relevant global information.
-This dual attention strategy enhances the model’s capacity to reconstruct high-resolution images with reduced computational demands.
-In parallel, the Deep Multilevel Feature Fusion Network (DMFFN) [15] addresses common challenges in super-resolution tasks, such as blurred edges and redundant structures.
-Using nested two-level residual blocks with asymmetric structures, DMFFN efficiently fuses multilevel features, leading to improved edge preservation and faster convergence during training.
-These contemporary models demonstrate significant progress in balancing performance and efficiency.
-Their architectures are designed to maintain structural consistency and fine details, addressing the limitations observed in earlier methods.
-Perceptual metrics such as LPIPS [16] have become widely adopted for evaluating visual similarity in a way that aligns closely with human perception.
-Beyond perceptual loss, incorporating visual cues that guide perception—namely color, brightness, contrast, and sharpness—is equally important.
-We select these attributes based on extensive literature: [17] identifies them as key appearance properties in image quality assessment; [18] highlights the role of color in quality estimation; [19] emphasizes the critical balance of brightness and contrast in enhancement; and [20] combines contrast and sharpness to assess enhancement reliability.
-By combining insights from generative models, reference-based super-resolution, attention mechanisms, and perceptual consistency techniques, we target high-quality, structurally accurate image synthesis suited for immersive visual applications.
+Recent work includes LatticeNet [12], which uses lattice-style blocks for adaptive feature fusion with low computation, enabling high-quality reconstruction without deep networks.
+However, its reliance on localized operations limits its capacity to capture long-range dependencies and global context, making it less effective in scenes with large structural variation or complex spatial arrangements.
+ELAN [13] combines shift convolutions with group-wise multiscale self-attention to capture both local and long-range dependencies, achieving a balance of accuracy and efficiency for real-time tasks.
+Nevertheless, the use of fixed attention windows introduces boundary discontinuities and restricts the model’s ability to fully capture cross-region interactions, particularly in non-uniform or geometrically diverse image regions.
+Further exploring attention-based models, DLGSANet [14] employs a hybrid of Multi-Head Dynamic Local Self-Attention (MHDLSA) for local features and Sparse Global Self-Attention (SparseGSA) for key global cues, improving high-resolution reconstruction with lower compute.
+While this hybrid structure improves reconstruction quality, it models local features implicitly and may introduce irrelevant global tokens, leading to suboptimal detail preservation in fine-grained textures or highly cluttered scenes.
+Similarly, DMFFN [15] employs nested two-level asymmetric residual blocks to fuse multilevel features, enhancing edge preservation and accelerating convergence.
+However, its fixed fusion design limits adaptability across spatial scales, and deeper residual nesting can introduce training instability.
+These architectures, while advancing performance-efficiency trade-offs, may still face challenges in maintaining structural consistency under complex geometric variations.
+To address these limitations, we propose a cross-view attention-based framework that explicitly models global context, aligns geometric structure across views, and adaptively fuses multi-scale features, leading to more robust reconstruction under challenging structural and visual conditions.
+By combining insights from generative models, reference-based super-resolution, attention mechanisms, and perceptual consistency techniques, our approach targets high-quality, structurally accurate image synthesis suited for immersive visual applications.
 3 Method
-The core philosophy of our solution is to use a cross-view attention mechanism to fuse content and visual quality features from a given pair of wide and narrow-angle images.
-The images are initially partitioned into fixed-size patches.
-We pick a pair of wide and narrow patches from the pool and pass them across the network architecture, as described below.
+We define the task of perceptually enriching wide-angle imagery as a guided resolution enhancement problem.
+The model receives a pair of co-captured images: a wide field-of-view (FoV) image that offers spatial context, and a narrow FoV image that retains fine-grained visual fidelity.
+The output is a physically upscaled version of the wide FoV image, enriched with perceptual cues derived from the narrow FoV capture.
+While the model performs a spatial upscaling operation on the wide FoV input, the true goal is not just to increase pixel count, but to enrich the image perceptually.
+The resolution increase provides a larger canvas, enabling the injection of meaningful details such as sharp textures, balanced contrast, and natural colors.
+A detailed, step-by-step walk-through of the methodology is provided below.
 3.1 Image Partitioning
-Both narrow and wide FoV images are partitioned into $128\times 128$-sized patches to ensure portions of images sharing visual similarity are used to exchange content and visual quality data.
+We start by dividing both the narrow and wide FoV images into fixed-sized patches to ensure portions of images sharing visual similarity are used to exchange content and visual quality data.
 Once partitioned, patches from wide and narrow FoV images that share visual properties are mapped using perceptual loss [16] and stored in the pool, ready to be pushed into the generator network.
 A cosine similarity on the feature maps generated by the mid layers of a VGG-19 [17] backbone is used to determine the perceptual distance between patches.
+During inference, as further explained in the experimentation section, the same partitioning ratio is maintained.
 Figure 2: Image partitioning and patch similarity between narrow and wide FoV images.
-Figure 3: Detailed architecture for the Generator component
 3.2 Network Architecture
 The architecture of the generator network is built upon a stack of ResNet-like [18] residual blocks that take as inputs a pair of wide and narrow FoV patches from the pool.
-The patches are encoded using separate encoder components to capture structural (content) and visual features from wide and narrow FoV patches respectively.
-Each block then employs a cross-view attention mechanism to fuse the content and visual properties from wide and narrow feature maps respectively.
-The pipeline consists of the following major components,
+The patches are encoded using separate encoder components to capture structural (content) and visual features from wide and narrow FoV patches, respectively.
+Each block then employs a cross-view attention mechanism to fuse the content and visual properties from wide and narrow feature maps, respectively.
+The pipeline consists of the following major components and has been visually illustrated in Fig 1.
 3.2.1 Structural Feature Encoder
 Wide FoV patches are encoded with traditional CNN-based encoders to capture the content details. Encoded wide patch represented as $E_{w}$
+Figure 3: Detailed architecture for the Generator component
 3.2.2 Visual Feature Encoder
-These encoders are specialized in extracting the visual parameters from a patch and are used to encode the narrow FoV patches.
-We use the Gram matrix approach to represent the visual features of a patch and include the following additional parameters in the Gram matrix representation: color, brightness, contrast, and sharpness.
-The feature map is constructed as follows,
-  • The original feature map $F_{l}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ is generated using a VGG-19 network.
-  • Color information for each channel over spatial dimensions, also referred to as global average pooling [24], is computed as:
-    $[F_{l_{\text{color}}}]_{c}$ $=\frac{1}{H_{l}W_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}F_{l}(c,h,w),$
-    for $\quad\text{for }c=1,\dots,C_{l}$
-  • Global mean pixel intensity across all channels and spatial dimensions indicates the perceived luminance (brightness) of the patch. Computed as:
-    $F_{l_{\text{brightness}}}=\frac{1}{C_{l}H_{l}W_{l}}\sum_{c=1}^{C_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}F_{l}(c,h,w)$
-  • We represent the overall contrast of the patch using the standard deviation:
-    $F_{l_{\text{contrast}}}=\sqrt{\frac{1}{C_{l}H_{l}W_{l}}\sum_{c=1}^{C_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}\left(F_{l}(c,h,w)-\mu_{l}\right)^{2}}$
-    where $\mu_{l}=F_{l_{\text{brightness}}}$ is the global mean defined above.
-  • Sharpness is accounted for using the Laplacian operated, as mentioned in this paper [25]:
-    $F_{l_{\text{sharpness}}}(c,h,w)=(F_{l}(c,\cdot,\cdot)*\nabla^{2})(h,w)$
-    where * denotes the convolution and $\nabla^{2}$ is the Laplacian kernel.
-  • Finally, we augment the original feature map obtained from VGG-19 as follows:
-    $\hat{F}_{l}=[F_{l},F_{l_{\text{color}}},F_{l_{\text{brightness}}},F_{l_{\text{contrast}}},F_{l_{\text{sharpness}}}]$
-    where scalar or vector features are broadcast to match the spatial dimensions of $F_{l}$.
-Once the enhanced feature map has been computed, we perform a dot product between the enhanced feature map and its transpose to attain the Gram matrix representation $E_{n}$, carrying the visual essence of the narrow FoV patch.
-$E_{n}=\hat{F}_{l}\hat{F}_{l}^{\top}$
+Narrow FoV patches are passed through encoders that are specially crafted to capture the visual essence of an image.
+We intend to do so by initially computing the Gram matrix representation of the patch and then augmenting it with visual cues computed from the patch.
+The augmented Gram matrix is then termed as $E_{n}$.
+As mentioned earlier, we take the color, brightness, contrast, and sharpness parameters of a patch as the representative of its perceptual indicators.
+Detailed mathematical formulations are included in the supplementary material(Section 1).
 3.2.3 Cross-Attention Residual Blocks
 Once the patches are encoded into the required latent space, we employ an attention-based fusion mechanism that fetches the most relevant visual features from the narrow patch and fuses those with the content features generated from the wide patch.
+This step facilitates the propagation of high-frequency details into the resultant image.
   • Encoded wide FoV patch $E_{w}$, representing the content data, acts as the query matrix
-  • Encoded narrow FoV patch $E_{n}$, carrying the visual properties, provides both the key and value matrices
-  • The attention scores, which gather semantically aligned, superior quality visual features present in the encoded narrow FoV patch, are computed using the scaled dot product equation introduced in the attention paper [26],
-    $Attention(E_{n},E_{w})=softmax(\frac{E_{w}E_{n}^{\top}}{\sqrt{d_{k}}})E_{n}$
-  • The gathered attention scores are then incorporated into the wide FoV feature map using a residual addition, producing the feature map representation of the enhanced wide FoV patch, $E$
-    $E=E_{w}+Attention(E_{n},E_{w})$
+  • Encoded narrow FoV patch $E_{n}$, carrying the visual properties, provides both the key and value matrices. $d_{n}$ is the dimension of the matrix $E_{n}$
+  • The attention scores, which gather semantically aligned, superior quality visual features present in the encoded narrow FoV patch, are computed in equation (1), using the scaled dot product equation introduced in the attention paper [19],
+    $Attention(E_{n},E_{w})=softmax(\frac{E_{w}E_{n}^{\top}}{\sqrt{d_{n}}})E_{n}$ (1)
+  • The gathered attention scores are then incorporated into the wide FoV feature map using a residual addition in equation (2), producing the feature map representation of the enhanced wide FoV patch, $E$
+    $E=E_{w}+Attention(E_{n},E_{w})$ (2)
 3.2.4 Upsampling Layers
-The output of the final residual block is passed through a pixel shuffle [27] layer with a scale factor equal to 2 to super-resolve the patch, denoted by $E^{*}$.
+The output of the final residual block is passed through a pixel shuffle [20] layer, spatially upscaling the patch to accommodate the incorporated perceptual cues, denoted by $E^{*}$.
+Generator architecture has been illustrated in Fig 3
 3.3 Loss Calculation
 Generator loss is calculated patch-wise, i.e., each of the generated patches is compared against both the original wide and narrow FoV patches to quantify the content and visual deviations from the ground truth.
-Then the patches are assembled using seamless cloning, to eliminate visible seams during blending, and the reconstructed image is passed through the discriminator which computes a seam consistency loss over the entire image.
+Then the patches are assembled using seamless cloning to eliminate visible seams during blending, and the reconstructed image is passed through the discriminator, which computes a seam consistency loss over the entire image.
 The losses are described as follows,
@@
     Content Loss: We find the deviations of the contents of the generated patch and ground truth with the following pixel-wise difference,
-    $L^{(l)}_{content}(E^{o}_{w},E^{g}_{w})=\frac{1}{n}\sum_{i}(E^{o}_{w}-E^{g}_{w})^{2}$
-    Here $E^{o}_{w}$ and $E^{g}_{w}$ are the latent space representations of the original and generated patches.
+    $L^{(l)}_{content}(E^{o}_{w},E^{g}_{w})=\frac{1}{n}\sum_{l}(E^{o}_{w}-E^{g}_{w})^{2}$ (3)
+    Here $E^{o}_{w}$ and $E^{g}_{w}$ are the latent space representations of the original and generated patches, and $n$ is the dimension.
   2. 2.
-    Visual Quality Loss: This loss is calculated as the difference of the Gram matrix representation of the generated patch and the narrow FoV patch using the following equation,
-    $L^{(l)}_{visual}=\sum_{l}w_{l}||E^{o}_{n}-E^{g}_{n}||^{2}$
+    Visual Quality Loss: This loss is calculated as the weighted sum of differences of the Gram matrix representations of the generated patch and the narrow FoV patch, from mid layers of the VGG-19 backbone, using the following equation,
+[…]
     Here $E^{g}_{n}$ and $E^{o}_{n}$ are the Gram matrix representations of the patches from the generated and the original narrow FoV patches, respectively.
+$w_{l}$ is the weight assigned to the layer $l$.
   3. 3.
     Total generator loss can be written as,
-    $L^{(l)}_{G}=L^{(l)}_{content}+L^{(l)}_{visual}$
+    $L^{(l)}_{G}=L^{(l)}_{content}+L^{(l)}_{visual}$ (5)
 3.3.2 Discriminator Loss
@@
     Seam Consistency Loss This loss is aimed at eliminating the seams between the patches in the synthetic image by minimizing the perceptual difference at the boundaries of the adjacent patches.
-For adjacent patches $P_{i}$ and $P_{j}$, we take neighboring regions $R_{i}$ and $R_{j}$ which are $w$ pixels wide and compute the loss as,
-    $L^{(l)}_{seam}=\frac{1}{w}\sum||\phi_{l}(R_{j})-\phi_{l}(R_{i})||^{2}$
+For adjacent patches $P_{i}$ and $P_{j}$, we take neighboring regions $R_{i}$ and $R_{j}$ which are $b$ pixels wide and compute the loss as,
+[…]
   2. 2.
-    Perceptual Loss against Ground Truth To ensure the essence of the narrow FoV has been plausibly transferred to the synthetic image, the output is perceptually compared to the original wide FoV ground truth.
+    Perceptual Loss against Ground Truth To ensure the essence of the narrow FoV has been plausibly incorporated into the synthetic image, the output is perceptually compared to the original wide FoV ground truth.
 We use the same VGG19 activation maps to find the perceptual difference $L^{(l)}_{perceptual}$ between the generated wide FoV image and the original image from the dataset.
     Total discriminator loss is defined as follows,
-    $L^{(l)}_{D}=L^{(l)}_{seam}+L^{(l)}_{perceptual}$
+    $L^{(l)}_{D}=L^{(l)}_{seam}+L^{(l)}_{perceptual}$ (7)
 3.4 Cascading Lens Stack
@@
 With 90,000 high-quality images of 1024 x 1024 resolution, it includes a diverse range of urban and natural environments.
-These high-quality landscape images with rich textures and details challenge the method’s ability to handle complex natural scenes.
-We have randomly sampled 20000 + 500 images (discussed in the Training Details section) and fed into the training loop in the ascending order of pixel variance [29], to ensure “easier” images are encountered first.
+These high-quality landscape images with rich textures and details provide a rigorous training ground for the model to learn how to handle complex natural scenes.
+We have randomly sampled 20000 + 500 images (discussed in the Training Details section) and fed them into the training loop in the ascending order of pixel variance [22], to ensure “easier” images are encountered first.
+Furthermore, to accommodate the rectangular aspect ratios of ultra-wide images encountered in real-world scenarios, we have fine-tuned the model on the DIV2K dataset [23], which consists of 800 images at 2K (1920 × 1080; 16:9) aspect ratios, thereby facilitating the model’s ability to generalize effectively across varying aspect ratios.
 Figure 4: Convergence of loss curves shown over multiple iterations
-[H] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 2x Zoom 1x Zoom 1x Output [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 1x Output 0.5x Zoom 0.5x Output
+[H] Narrow Input Wide Input Output [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 2x Zoom 1x Zoom 1x Output [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 1x Output 0.5x Zoom 0.5x Output
 Table 1: Experimental results demonstrating the cascading effect of multi-zoom image upscaling.
@@
 This intentional degradation process mimics the wide image of low-quality imaging conditions while preserving the broader scene context, allowing us to test our methods under challenging visual conditions.
-4.3 Generator
-The generator takes in a pair of patches, each of dimension $128\times 128\times 3$.
-The wide FoV patch is passed through a CNN encoder with a final batchNorm layer to produce the content feature map.
-The narrow FoV patch initially undergoes feature extraction through the $Conv2_{2},Conv3_{2},Conv4_{2}$ layers of a VGG-19 backbone.
-The feature map produced by VGG-19 is enhanced by the visual properties computed from the original patch.
-We use 10 attention-based residual blocks, each with two $3\times 3$ convolution units, a ReLU, followed by the cross-view attention score calculation unit, outputs of which are added to the wide FoV patch.
-The details-infused wide FoV patch is then passed through a series of convolution layers interleaved by pixel shuffle layers to physically upscale the patch to a dimension of $256\times 256\times 3$.
-4.4 Discriminator
-We use a ResNet18 [31] based discriminator network that takes as input a $1024\times 1024\times 3$ image formed by stitching the enhanced patches generated by the generator using Poisson Blending [32].
-The image undergoes a series of $6$ residual blocks followed by a layer of adaptive pooling and a fully connected layer outputting a single scalar denoting the realness of the input image.
-4.5 Enhanced image reconstruction pipeline
-Figure 5 highlights the steps undertaken to reconstruct the enhanced.
+4.3 Image Reconstruction Pipeline
+During training, each image is passed through a reconstruction pipeline.
+The cycle begins with loading a pair of images, wide and narrow FoV, from the dataset, applying the preprocessing mentioned earlier, and dividing the processed images into patches.
+The patches, from the pool, are then fed into the generator, which produces their enhanced counterparts.
+The enhanced patches are stitched back into a single image, the super-resolved image, and forwarded to the discriminator for loss computation.
+This pipeline has been illustrated in Fig 5.
+Implementation details about the generator and discriminator components are provided in the supplementary document (Section 2).
 Figure 5: Enhanced image reconstruction pipeline
+Figure 6: Visual comparison of our outcomes with existing methods.
+Dataset ESRGAN LatticeNet ELAN DLGSANet DMFFN Ours
+RealSRv3 0.2940 0.2907 0.2582 0.3409 0.3054 0.1740
+ImagePairs 0.2979 0.2775 0.2754 0.2866 0.2872 0.2435
+Table 2: LPIPS Comparison (best marked in red)
+Dataset ESRGAN LatticeNet ELAN DLGSANet DMFFN Ours
+RealSRv3 32.39 32.82 32.69 32.79 33.13 33.61
+ImagePairs 29.21 29.23 29.13 29.09 29.09 29.29
+Table 3: PSNR Comparison (highest marked in red)
+Dataset ESRGAN LatticeNet ELAN DLGSANet DMFFN Ours
+RealSRv3 0.8385 0.8479 0.8582 0.8433 0.8557 0.8870
+ImagePairs 0.7244 0.7168 0.6782 0.7421 0.7604 0.7597
+Table 4: SSIM Comparison (highest marked in red)
 4.4 Training Details
-Key hyperparameters for training and evaluation included a patch size of 128 × 128, Adam optimizer with a learning rate of 0.0001 with exponential decay every 10 epochs, a batch size of 16, perceptual loss layers (Conv2_2, Conv3_2, Conv4_2 of VGG19), and a threshold for patch similarity empirically set to 0.7.
+Key hyperparameters for training and evaluation included an Adam optimizer with a learning rate of 0.0001 with exponential decay every 10 epochs, a batch size of 16, perceptual loss layers (Conv2_2, Conv3_2, Conv4_2 of VGG19), and a threshold for patch similarity empirically set to 0.7.
+During training, an image is partitioned into 64 patches, and the patch dimensions are determined likewise, for example $128\times 128$ for Landscape dataset images.
+In case of inference, image dimensions are resized to the next multiple of 64, if needed, to facilitate the partitioning strategy.
 The generator network was pre-trained for 50 epochs using content and visual losses before incorporating adversarial loss, each epoch carrying 10 sample images.
 The full generator-discriminator framework was then trained for an additional 2000 epochs, totaling 20.5K iterations.
-The convergence behavior of the generator loss has been illustrated in Figure 4.
+The convergence behavior of the generator loss has been illustrated in Fig 4.
 These curves represent the first 15 epochs where the generator has been trained solo.
 Upper row features the raw loss curves as generated by the GAN architecture, while the lower row has been smoothened using the numpy convolve [25] method.
-Narrow FoV Input wide FoV Our result
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Table 2: Experimental results showing the output generated by our model when processing narrow and wide input images.
-[H] Dataset bicubic Multiscale GAN Ours Set5 0.5657 0.2823 0.0795 Set14 0.6114 0.3717 0.2032 BSD100 0.5726 0.4112 0.3361
-Table 3: LPIPS Comparison (best marked in red)
-Figure 6: Visual comparison of our results against the mentioned prior arts
-Dataset ESRGAN LatticeNet ELAN DLGSANet DMFFN Ours
-Set5 32.73 32.18 32.43 32.46 32.41 32.80
-Set14 28.99 28.61 28.78 28.79 28.74 30.37
-BSD100 25.16 27.57 27.69 27.70 27.74 26.58
-Table 4: PSNR Comparison (highest marked in red)
-Dataset ESRGAN LatticeNet ELAN DLGSANet DMFFN Ours
-Set5 0.9011 0.8943 0.8975 0.8993 0.8981 0.9118
-Set14 0.7917 0.7812 0.7858 0.7871 0.7874 0.7806
-BSD100 0.7455 0.7355 0.7406 0.7415 0.7421 0.8261
-Table 5: SSIM Comparison (highest marked in red)
+4.5 Experimental Setting
+For evaluation against existing works in the field, we have used 2 benchmark datasets which are RealSRv3 [26] and ImagePairs [27].
+These evaluation datasets differ from classical SR datasets in that the narrow and wide-FoV pairs are captured using actual focal length manipulation, as opposed to mathematical simulation, and hence are better suited to our particular case.
+We have, however, performed a parallel validation on traditional datasets Set5 [28], Set14 [29], and BSD100 [30], results of which are available in the supplementary document (Section 4).
+All reported metrics have been calculated based on the complete output image, as opposed to patch-wise measures.
+We have used the LR image provided by the dataset as the input, and the given HR as the ground truth.
 4.6 Experiment Results
-Table 2 presents the results of our experiment, where a narrow and a low-resolution wide image are used as input. These images are processed through our model to generate the final output.
 Table 1 illustrates the cascading effect of the outputs derived from images captured at different zoom levels.
@@
 This sequence effectively simulates how our method can be applied in a cascading manner using all available zoom lenses.
+Visual outcomes on the benchmark datasets have been elaborated in the supplementary document (Section 3B).
+Loss convergence in the epoch interval of 5 to 10 has been shown in Fig 4.
+The lower curve has been smoothed to denote the overall downward trend of the graph.
+Details on loss minimization has been included in the supplementary document (Section 3A).
 5 Comparison with Prior Arts
-We have evaluated our proposed methods on PSNR, SSIM and LPIPS [16] metrics over Set5 [34], Set14 [35] and BSD100 [36] datasets.
-For PSNR and SSIM, we have compared our results with existing technologies such as ESRGAN [7], LatticeNet [12], ELAN [13], DLGSANet [14] and DMFFN [15] in Table 4 and Table 5.
-We have presented our comparison with bicubic interpolation and Multiscale GAN [37] on the LPIPS metric in Table 3.
-Visual comparison of our method against prior super-resolution models has been presented in Figure 6.
+We have evaluated our proposed methods on LPIPS [31], PSNR, and SSIM metrics over datasets mentioned earlier.
+For all the metrics, we have compared our results with existing technologies such as ESRGAN [7], LatticeNet [12], ELAN [13], DLGSANet [14] and DMFFN [15] and presented the results in Tables 4, 4, and 4 respectively.
+We have also included a visual comparison against the prior methods in the Fig 6.
+5.1 Quantitative Analysis
+In terms of perceptual quality, as assessed using the LPIPS metric, our approach shows a clear improvement over earlier methods.
+This is largely due to the integration of visually similar patches, the selection of representative perceptual features, and the inclusion of perceptual loss within the discriminator, all of which contribute to more visually realistic outputs.
+Similarly, our model achieves a noticeably higher PSNR score compared to other techniques. Notably, on the ImagePairs dataset, it is the only method to surpass a PSNR value of 30.
+In the case of SSIM, our method also achieves a marked advantage on the RealSRv3 dataset, indicating superior preservation of structural details.
+However, for the ImagePairs dataset, the DMFFN model slightly outperforms ours.
+This can likely be attributed to its Multi-stage Feature Supplementation, which enhances geometrical consistency by combining shallow and deep features—albeit with increased computational demands.
+5.2 Qualitative Analysis
+We evaluated 2× super-resolved outputs from the referenced datasets against those produced by existing models.
+First, many of these models lack broader content-aware guidance, making it difficult for them to accurately reconstruct structural details—this is especially evident in elements like building windows (first image).
+Second, earlier methods tend to produce less distinct edges, often resulting in soft or blurred boundaries, whereas our approach delivers noticeably sharper edges, as illustrated by the leaf contours in the second image.
+Lastly, in visually rich scenes with high color saturation and contrast—such as green grass (third image)—previous models often fail to preserve vibrancy and fine texture.
+In contrast, our method successfully maintains these visual attributes by utilizing cues from the narrow field-of-view input, which is rich in detail.
 6 Conclusion
-In this work, we presented a novel approach for enhancing ultra-wide field-of-view (FoV) images by leveraging paired narrow-FoV views of the same scene.
-The core of our method is a patch-based cross-view attention mechanism that selectively transfers fine-grained details from narrow-FoV patches to their wide-FoV counterparts, guided by Gram matrix-based visual encoding that captures texture statistics for accurate alignment.
+This paper introduces an approach for enhancing ultra-wide field-of-view (FoV) images by leveraging paired narrow-FoV views of the same scene.
+The core of our method is a patch-based cross-view attention mechanism that selectively synthesizes fine-grained details from narrow-FoV patches into their wide-FoV counterparts, guided by a Gram matrix-based visual encoding to capture the texture statistics for accurate alignment.
 To further refine the reconstruction, we introduce a cascading lens stack mechanism that progressively transfers details across adjacent lens images, facilitating spatial consistency and coherent integration across the full image.
-We trained our model on the Landscape HQ dataset to ensure coverage of both structural diversity and fine detail.
-For evaluation, we used standard benchmark datasets including BSD100, Set5, and Set14.
-Experimental results demonstrate that our method consistently outperforms strong baselines such as ESRGAN, ELAN, DLGSANet, DMFFN, and LatticeNet in both PSNR and SSIM metrics.
-Qualitative comparisons further confirm the perceptual superiority of our results.
-Overall, our framework successfully addresses the challenge of enhancing ultra-wide FoV images by combining cross-view attention, texture-aware encoding, and a cascading lens-based reconstruction strategy.
-Future extensions could explore its application to video enhancement and real-time inference scenarios.
+The model trained is on the Landscape HQ dataset to cover both structural diversity and finer details, and fine-tuned on DIV2K for handling rectangular images.
+For evaluation, we have used standard benchmark datasets including RealSRv3 and ImagePairs.
+Experimental results demonstrate a promising improvement over strong baselines such as ELAN, DLGSANet, DMFFN, and LatticeNet, especially in LPIPS, along with both PSNR and SSIM metrics.
+Qualitative comparisons further add to the perceptual superiority of our results.
+Limitations to our current work include the necessity of further fine-tuning our model with real-life ultra-wide and narrow FoV pairs, since it is trained on synthetic data.
+Also, additional optimizations are required to deploy the model on-device to serve real-time purposes and video super-resolution tasks.
+Declarations
+  • Funding: This research did not receive external funding.
+  • Competing Interests: The authors do not have any competing interests as defined by Springer, or other interests that might be perceived to influence the results and/or discussion reported in this paper.